OpenStudy (anonymous):

how long would it take for you to triple your money if you invest at a rate of 6% a year compounded annually?

OpenStudy (anonymous):

\[(1.06)^t=3\] solve for \(t\)

OpenStudy (anonymous):

in one step using \[b^x=A\iff x=\frac{\ln(A)}{\ln(b)}\]

OpenStudy (anonymous):

19 years 1.06^19=3.0256